# Beyond Ten Turns: Unlocking Long-Horizon Agentic Search with Large-Scale Asynchronous RL

### v3

@@ 4.3 Training Dynamics
In Fig. 9 and Fig. 10, we plot the number of generated tokens, search queries and webpage browsing for ASearcher-Local-7B and ASearcher-Local-14B training, respectively. With our training recipe, length increment and the increase in the number of tool callings are observed in both 7B and 14B scales. Notably, the number of search queries scale up to 6, which is higher than the numbers reported by prior works (Jin et al., 2025; Song et al., 2025). Interestingly, we find that the 7B model fails to learn valid webpage browsing while the 14B model can learn to access webpage to solve challenging questions in the later stage of training. We hypothesize that the failure of 7B model in learning webpage browsing occurs because the model capacity is too small to stably learning summarize lengthy webpages in a zero RL training setting.
@@ 5 Related Works
Some works have constructed agent workflows that enable large language models (LLMs) to leverage external tools for solving complex tasks, with notable examples include Search-o1Li et al. (2025c) and ReAgentZhao et al. (2025). Prompt-based methods, while effective for rapid development, are fundamentally limited by the capacity of the underlying LLMs and could not be reliably improved with environment feedback. Some works attempt to construct SFT trajectories for LLMs. For instance, Asai et al. (2023); Yu et al. (2024) leverage large LLMs to synthesize retrieval and reasoning trajectories to fine-tune smaller models. Recently, some works investigate Reinforcement learning (RL) methods to enhance the LLM-based agents, mostly focusing on multi-hop QA benchmarks such as HotpotQA and 2WikiMultihop. Jin et al. (2025); Song et al. (2025); Chen et al. (2025); Zheng et al. (2025) perform RL training with multi-hop QA data and observe an increase in the number of tool uses. RAG-R1 (Tan et al., 2025) further combines SFT and RL to enhance the search strategies. More recently, researchers have begun to focus on more challenging tasks, by fine-tuning sophisticated prompt-based agents powered by Large Reasoning Models through offline RL (Li et al., 2025d), SFT on simulated trajectories with real-world web data (Sun* et al., 2025; Li et al., 2025b), and constructing challenging QAs for RL training. (Tao et al., 2025).
@@
Rather than relying solely on large-scale human annotation, data synthesis has emerged as a scalable approach to prepare training data for search agents. Recent approaches generate synthetic but realistic QA trajectories by interacting with real web pages and curating data using LRMs Sun* et al. (2025); Wu et al. (2025a); Li et al. (2025b). On the other hand, WebSailor (Li et al., 2025b) constructs structurally challenging tasks through sampling and fuzzing, and WebShaper (Tao et al., 2025) utilizes techniques from set theory to construct high-quality complex QAs. By contrast, ASearcher develops an autonomous LLM agent for synthesizing challenging QAs with high uncertainty, without relying on complex knowledge graphs. Both the data synthesis agent and the synthetic training data in ASearcher are fully open-sourced.
@@ Appendix A Full Case Study
In this section, we provide a detailed case study on an extremely challenging question from GAIA (Mialon et al., 2023). Specifically, we analyze Search-R1-32B (Jin et al., 2025) and Search-o1 (QwQ) (Li et al., 2025c) in Fig. 11.

### v4

@@ 4.3 Training Dynamics
In Fig. 9 and Fig. 10, we plot the number of generated tokens, search queries and webpage browsing for ASearcher-Local-7B and ASearcher-Local-14B training, respectively. With our training recipe, length increment and the increase in the number of tool callings are observed in both 7B and 14B scales. Notably, the number of search queries scale up to 6, which is higher than the numbers reported by prior works [11, 30]. Interestingly, we find that the 7B model fails to learn valid webpage browsing while the 14B model can learn to access webpage to solve challenging questions in the later stage of training. We hypothesize that the failure of 7B model in learning webpage browsing occurs because the model capacity is too small to stably learning summarize lengthy webpages in a zero RL training setting.
@@
4.4 Emergent Behaviors
Figure 11: Left: Word count of reflective keywords during training time. Right: Word count of keywords indicating explicit reference of external information.
Keyword Analysis.
In Fig. 11, we plot the word count of different keywords during the training process of ASearcher-Web-QwQ. These keywords include words that indicate reflection behaviors of the agent, such as “alternatively”, “however”, and “wait”, and also words indicating that the agent is referencing to external information, such as “doc” and “mention”. From Fig. 11, it is clear that the agent learns to reflect over previous actions and conclusions. Interestingly, we also see that, in the second stage of RL training where the training focuses on challenging tasks, the agent learns to refer to external information, as evidenced by the increment in the word count of “doc” and “mention” after step 200.
@@ 5 Related Works
Some works have constructed agent workflows that enable large language models (LLMs) to leverage external tools for solving complex tasks, with notable examples include Search-o1[18] and ReAgent[48]. Prompt-based methods, while effective for rapid development, are fundamentally limited by the capacity of the underlying LLMs and could not be reliably improved with environment feedback. Some works attempt to construct SFT trajectories for LLMs. For instance, [4, 47] leverage large LLMs to synthesize retrieval and reasoning trajectories to fine-tune smaller models. Recently, some works investigate Reinforcement learning (RL) methods to enhance the LLM-based agents, mostly focusing on multi-hop QA benchmarks such as HotpotQA and 2WikiMultihop. [11, 30, 5, 49] perform RL training with multi-hop QA data and observe an increase in the number of tool uses. RAG-R1 [33] further combines SFT and RL to enhance the search strategies. More recently, researchers have begun to focus on more challenging tasks, by fine-tuning sophisticated prompt-based agents powered by Large Reasoning Models through offline RL [19], SFT on simulated trajectories with real-world web data [32, 17], and constructing challenging QAs for RL training. [34].
@@
Rather than relying solely on large-scale human annotation, data synthesis has emerged as a scalable approach to prepare training data for search agents. Recent approaches generate synthetic but realistic QA trajectories by interacting with real web pages and curating data using LRMs [32, 39, 17]. On the other hand, WebSailor [17] constructs structurally challenging tasks through sampling and fuzzing, and WebShaper [34] utilizes techniques from set theory to construct high-quality complex QAs. By contrast, ASearcher develops an autonomous LLM agent for synthesizing challenging QAs with high uncertainty, without relying on complex knowledge graphs. Both the data synthesis agent and the synthetic training data in ASearcher are fully open-sourced.
@@ Appendix A Full Case Study
In this section, we provide a detailed case study on an extremely challenging question from GAIA [24]. Specifically, we analyze Search-R1-32B [11] and Search-o1 (QwQ) [18] in Fig. 12.